# Some Plancherel identities for unbounded subsets of $\mathbb{R}$ in duality

### v1

@@ 1. Introduction
Let $\Omega$ be an open subset in $\mathbb{R}$ and let $p\in\mathbb{N}$, $p\geq 2$. Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
@@
Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$ and let $p\in\mathbb{N}$. We say that $E$ and $F$ are congruent $\operatorname{mod}p\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
@@
With Theorem 1.6, we get that $\Omega$ has spectrum $\mathbb{Z}$. Also, let $\Omega_{0,k}=\Omega_{0}\cap[k,k+1)$ and let $I_{k}=\Omega_{0,k}-k\subset[0,1)$, for all $k\in\mathbb{Z}$. Then the sets $\{I_{k}:k\in\mathbb{Z}\}$ are disjoint (otherwise the sets $\Omega_{0}+k$ are not), and their union is $[0,1)$ because, for $x\in[0,1)$ there is a unique $k\in\mathbb{Z}$ and $x_{0}\in\Omega_{0}$ such that $x=x_{0}-k$. So $x_{0}\in\Omega_{0,k}$ and $x\in I_{k}$. Thus $\Omega_{0}$ is congruent to $[0,1)$ modulo $\mathbb{Z}$ and
@@
(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2}.$
@@
(2.5) $\hat{\chi}_{\Omega}=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
@@
$f=\chi_{\Omega}\cdot\varphi.$
@@
$\hat{f}(t)=\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
@@
Let $a_{1}=-\frac{1}{2p}-b-\epsilon$, $a_{2}=-\frac{1}{2p}-b$. We have
@@
Consider two sets $\Omega_{0}+pl+j$ and $\Omega_{0}+pl^{\prime}+j^{\prime}$ with $l,l^{\prime}\in\mathbb{Z}$ and $j,j^{\prime}\in\{0,\dots,p-1\}$. If $j\neq j^{\prime}$ then the intersection of these two sets in contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$. If $j=j^{\prime}$ and the intersection of the two sets is nonempty, then $(\Omega_{0}+pl)\cap(\Omega_{0}+pl^{\prime})\neq\emptyset$ and this is impossible because $\Omega_{0}$ is contained in $[0,p)$. This proves (i).
@@
$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $[0,1)$.
For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and

### v2

@@ 1. Introduction
Let $\Omega$ be an open subset in $\mathbb{R}$ and let $p\in\mathbb{N}$, $p\geq 2$. Then $\Omega$ tiles $\mathbb{R}$ by $\{0,1,\dots,p-1\}$ if and only if $\Omega$ is a spectral set with pair measure $\mu$ equal to a renormalized Lebesgue measure on $\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}$, $\mu=p\operatorname*{Leb}_{\left[-\frac{1}{2p},\frac{1}{2p}\right]+\mathbb{Z}}$ .
@@
Let $E$ and $F$ be two measurable subsets of $\mathbb{R}$. We say that $E$ and $F$ are congruent $\operatorname{mod}\mathbb{Z}$ if there exists a partition (up to measure zero) of $E$ into possibly empty sets $\{E_{k}:k\in\mathbb{Z}\}$ such that $\{E_{k}+k:k\in\mathbb{Z}\}$ is a partition of $F$ (up to measure zero).
@@
With Theorem 1.6, we get that $\Omega_{0}$ has spectrum $\mathbb{Z}$. Also, let $\Omega_{0,k}=\Omega_{0}\cap[k,k+1)$ and let $I_{k}=\Omega_{0,k}-k\subset[0,1)$, for all $k\in\mathbb{Z}$. Then the sets $\{I_{k}:k\in\mathbb{Z}\}$ are disjoint (otherwise the sets $\Omega_{0}+k$ are not), and their union is $[0,1)$ because, for $x\in[0,1)$ there is a unique $k\in\mathbb{Z}$ and $x_{0}\in\Omega_{0}$ such that $x=x_{0}-k$. So $x_{0}\in\Omega_{0,k}$ and $x\in I_{k}$. Thus $\Omega_{0}$ is congruent to $[0,1)$ modulo $\mathbb{Z}$ and
@@
(2.1) $\sum_{\lambda\in\Lambda}|\hat{\mu}(t+\lambda)|^{2}=m^{2},$
where $\hat{\mu}$ is the Fourier transform of the measure $\mu$,
$\hat{\mu}(t)=\int e^{-2\pi itx}\,d\mu(x),\quad(t\in\mathbb{R}).$
@@
(2.5) $\hat{\chi}_{\Omega}=\frac{1}{p}\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}.$
@@
$f=p\chi_{\Omega}\cdot\varphi.$
@@
$\hat{f}(t)=p\hat{\chi}_{\Omega}\ast\hat{\varphi}(t)=\left(\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)\delta_{k/p}\right)\ast g(t)=\sum_{k\in\mathbb{Z}}\hat{\chi}_{\Omega_{0}}(k/p)g(t-k/p).$
@@
Let $a_{1}=\frac{1}{2p}+b+\epsilon$, $a_{2}=\frac{1}{2p}+b$, for $b\in\mathbb{R}$ and $\epsilon>0$. We have
@@
Consider two sets $\Omega_{0}+pl+j$ and $\Omega_{0}+pl^{\prime}+j^{\prime}$ with $l,l^{\prime}\in\mathbb{Z}$ and $j,j^{\prime}\in\{0,\dots,p-1\}$. If $j\neq j^{\prime}$ then the intersection of these two sets is contained in $(\Omega^{\prime}+j)\cap(\Omega^{\prime}+j^{\prime})=\emptyset$. If $j=j^{\prime}$ and the intersection of the two sets is nonempty, then $(\Omega_{0}+pl)\cap(\Omega_{0}+pl^{\prime})\neq\emptyset$ and this is impossible because $\Omega_{0}$ is contained in $[0,p)$. This proves (i).
@@
$(\Omega_{0}\cup R)+\mathbb{Z}=(\Omega_{0}+\mathbb{Z})\cup(R+\mathbb{Z})=((\Omega_{0}+\mathbb{Z})\cup R)+\mathbb{Z}\supseteq[0,1)+\mathbb{Z}=\mathbb{R}.$
This shows that $\Omega_{0}\cup R$ tiles $\mathbb{R}$ by $\mathbb{Z}$.
For $\tilde{\Omega}=(\Omega_{0}\cup R)+p\mathbb{Z}$, since $\Omega_{0}\cup R$ tiles $\mathbb{R}$ with $\mathbb{Z}$, we get that the sets $\{\tilde{\Omega}+j=(\Omega_{0}\cup R)+j+p\mathbb{Z}:j\in\{0,\dots,p-1\}\}$ are disjoint, and
@@
Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data Availability Statement. No new data were created or analyzed in this study. Data sharing is not applicable to this article.
Funding. No funds, grants, or other support were received during the preparation of this manuscript.
Author Contribution. All authors contributed equally to the study conception, design, data collection, analysis, and manuscript preparation. Both authors read and approved the final manuscript.